# Nonparametric Least squares estimators for interval censoring

### v4

@@ Abstract
The limit distribution of the nonparametric maximum likelihood estimator for interval censored data with more than one observation time per unobservable observation, is still unknown in general. For the so-called separated case, where one has observation times which are at a distance larger than a fixed $\epsilon>0$, the limit distribution was derived in [piet:96]. For the non-separated case there is a conjectured limit distribution, given in [GrWe:92], Section 5.2 of Part 2. But the findings of the present paper suggest that this conjecture may not hold.
We prove consistency of a related nonparametric isotonic least squares estimator and give a sketch of the proof for a result on its limit distribution. We also provide simulation results to show how the nonparametric MLE and least squares estimator behave in comparison. Moreover, we discuss a simpler least squares estimator that can be computed in one step, but is inferior to the other least squares estimator, since it does not use all information.
For the simplest model of interval censoring, the current status model, the nonparametric maximum likelihood and least squares estimators are the same. This equivalence breaks down if there are more observation times per unobservable observation. The computations for the simulation of the more complicated interval censoring model were performed by using the iterative convex minorant algorithm. They are provided in the GitHub repository [piet_intcens:25].
@@ 1 Introduction
The simplest and most studied interval censoring model is the so-called interval censoring, case 1, or current status model. This model can be defined in the following way (see, e.g., Section 2.3 of [piet_geurt:14]).
@@
Since distribution functions are by definition nondecreasing, computing the maximum likelihood estimator poses a shape constrained optimization problem. As can be seen from (1.2), the value of $\ell$ only depends on the values that $F$ takes at the observed time points $T_{i}$. Hence one can choose to consider only distribution functions that are constant between successive observed time points $T_{i}$. The lemma below shows that this estimator can be characterized in terms of a greatest convex minorant of a certain diagram of points. The following result is Lemma 2.7 in [piet_geurt:14] (and Proposition 1.2 in [GrWe:92]).
@@
[Characterization of the nonparametric ML estimator in the current status model] Consider the cumulative sum diagram consisting of the points $P_{0}=(0,0)$ and
@@
over all monotone nondecreasing (not necessarily bounded by 0 and $1$) functions $F$, see p. 43 in section 1.1 of part 2 of [GrWe:92]. It also follows from Theorem 1.2.1 on p. 7 of [rwd:88], where the connection brtween the derivative of the greatest convex minorant and the solution of the isotonic regression problem is given.
So we have the two characterizations of the nonparametric maximum likelihood estimator, the characterization as a maximizer of (1.2) and the characterization as a minimizer of (1.3) (see, e.g., [GrWe:92]). Note that the weights are equal and constant in the least sauares problem.
@@
where $\Delta_{i2}=1-\Delta_{i0}-\Delta_{i1}$. It is not clear that there is an equivalent to the minimization of (1.3) in this situation. We cannot use Theorem 1.5.1 in [rwd:88] because we have the difference of $F(V_{i})$ and $F(U_{i})$ in the terms $\log(F(V_{i})-F(U_{i}))$ instead of just $F(U_{i})$ or $F(V_{i})$ by itself only.
@@
Figure 2: (a) Simulated variances, times $n^{2/3}$, of the nonparametric MLE (black solid curve) and the least squares estimate (red) minimizing (1.6), for $t_{i}=0.1,0.2,\dots,1.9$, linearly interpolated between values at the $t_{i}$ for the model of Example 1. The blue dashed curve is the theoretical limit curve one obtains from Theorem 1 in Section 2 for the LS estimator. The simulated variances are based on $10,000$ simulations of samples of size $n=1000$ for the truncated exponential distribution function $F_{0}$ on $[0,2]$ and the order statistics of the uniform distribution on $0,2]^{2}$ as observation times. (b) The same comparison, but now for $F_{0}$ uniform on $[0,2]$.
@@
Inspired by the conjectured faster rate of convergence of the MLE in the non-separated case (i.e., observation intervals can be arbitrarily small) in Section 5.2 of Part 2 of [GrWe:92], Lucien Birgé constructed in [lucien:99] a histogram estimator which actually achieves rate $(n\log n)^{1/3}$ locally (but suffers severely from bias, in contrast with the MLE). The asymptotic (normal) distribution of the histogram estimator was derived in [piet_tom:11]. At the time that [piet_tom:11] was written, it was still believed that the nonparametric MLE would also achieve this faster rate. The matter has still not been settled, but the present paper casts doubts on this conjecture.
It is also possible to define a least squares estimator which can be computed in one step, using the convex minorant (or “pool adjacent violators”) algorithm. An estimator of this type was proposed for the uniform deconvolution problem in [eszuijlen:96]. This estimator minimizes
@@
We compare the two least squares estimators for sample size $n=10,000$ in Figure 3, from which the superiority of the estimator, based on minimizing (1.6), seems clear.
@@
Our proof of the limit result Theorem 1 is not complete. We still have to show that the “off-diagonal terms” (5.8) and (5.9) are $o_{p}(n^{-2/3})$, see Section 5.2. In fact, terms of this type were shown to be $O_{p}(n^{-5/6})$ in a similar problem, analyzed in [piet:24]. In Section 3 we discuss the theory, needed for the completion of the proof.
@@ 2 The least squares estimators
The proof is based on the so-called Fenchel duality condition, as also used in [piet_geurt:14] in the characterization of the nonparametric MLE for interval censoring (apart from the Lagrange multipliers in the present case) and is therefore omitted. A picture of the process $W_{n,\hat{F}_{n}}$ for $n=100$ is given in Figure 4 for Example 1 above. $W_{n,\hat{F}_{n}}$ touches zero at points just to the left of points where the least squares estimator $\hat{F}_{n}$ has a jump. The iterative convex minorant algorithm for computing $\hat{F}_{n}$ is based on this characterization.
@@
Lemma 3.
Let $\hat{F}_{n}$ be the isotonic least squares estimator, minimizing (1.6) over distribution functions $F$ under the conditions of Theorem 1 below. Then $\hat{F}_{n}$ converges almost surely to $F_{0}$ in the supremum metric.
Proof.
Let the function $\psi$ be defined by
$\psi(F)$
Since $\hat{F}_{n}$ minimizes (2) over $F$, we must have
$\geq 0.$
Proceeding as in Section 4 of Part 2 of [GrWe:92], using the Helly compactness theorem, we get from this, for a limit point $F$ of a subsequence of $\hat{F}_{n}$:
$\int\{F_{0}(u)-F(u)\}\{F(u)-F_{0}(u)\}\,h(u,v)\,du\,dv$
$\qquad+\int\{F_{0}(v)-F_{0}(u)-F_{(}v)-F(v)+F(u)\}$
$\qquad+\int\{F_{0}(v)-F(v)\}\{F(v)-F_{0}(v)\}\,h(u,v)\,du\,dv$
$\qquad-\int\{F(v)-F_{0}(v)\}^{2}\,h(u,v)\,du\,dv$
$\geq 0.$
@@
Let $F_{0}$ have a continuous positive density $f_{0}$ on $[0,M]$ and let $(U,V)$ be the order statistics of an absolutely continuous distribution. We assume that $(U,V)$ has a a positive continuous density $h$ on
@@
with first and second marginals $h_{1}$ and $h_{2}$, respectively, with bounded second partial derivatives.We assume that $X_{i}$ is independent of $(U_{i},V_{i})$.
@@
$\sigma_{t_{0}}=\left(a_{t_{0}}f_{0}(t_{0})/b_{t_{0}}\right)^{1/3},\qquad t_{0}\in(0,M).$ (2.7)
@@
The conditions on the observation density $h$ are similar to the conditions in [GeGr:99] and used in the smooth functional theory needed in treating so-called “off-diagonal terms”, see Section 5.2.
@@
As an example, the asymptotic variance of the LS estimator is shown in Figure 5 as the blue dashed curve for in Example 1. Remarkably, the curve of the variances of the MLE is even closer to the theoretical asymptotic curve than the curve for the variances of the least squares estimator for this sampe size of $n=10,000$. This should be compared with Figure 2, where the sample size was $n=1000$ and the variances for the LS estimator were actually smaller that those of the nonparametric MLE. The faster decrease of the variances of the MLE than those of the LS when we increase the sample size from $1000$ to $10,000$ might give a hint of the possibility of a rate of $(n\log n)^{1/3}$ again, but the form of the curve is certainly not in accordance with the conjecture in [GrWe:92].
We here make some additional remarks about the proof of Theorem 1, discussed in the Appendix. Results of this type were proved for the nonparametric MLE in [piet:96] and [piet:24]. One has to show that the leading asymptotic behavior is provided by replacing the estimator by the underlying distribution function $F_{0}$ in the characterization, given in Lemma 2. To this end, one has to show that certain terms, called the “off-diagonal terms”, are of lower order. Doing this is rather hard in our experience ([piet:24] needed 53 pages), but presently there seems to be no other way. We give a sketch of these calculations in Section 3 and in the Appendix. For similar calculations, see, e.g., Chapter 10 of [piet_geurt:14] and [piet:24]. The first result of the type of Theorem 1 was for the current status model in Theorem 5.5 of Part 2 of [GrWe:92].
@@
So, in constructing the cusum diagram, we run through all $2n$ observation points; if we meet a point $U_{i}$, we add $\delta_{i0}$ to the cusum diagram, if we meet a point $V_{i}$ we add $\delta_{i0}+\delta_{i1}$ to the diagram. In the latter case we only record whether the unobservable variable lies to the left of $V_{i}$ (then $\delta_{i0}+\delta_{i1}=1$) or to the right of $V_{i}$ (then $\delta_{i0}+\delta_{i1}=0$).
@@
$(a_{t_{0}}^{\prime})^{2}$ $=F_{0}(t_{0})\{1-F_{0}(t_{0})\}\bigl\{h_{1}(t_{0})+h_{2}(t_{0})\bigr\}$ (2.10)
@@
$b_{t_{0}}^{\prime}$ $=\left\{h_{1}(t_{0})+h_{2}(t_{0})\right\}/4,$ (2.11)
@@
$\sigma_{t_{0}}^{\prime}=\left(a_{t_{0}}^{\prime}f_{0}(t_{0})/b_{t_{0}}^{\prime}\right)^{1/3},\qquad t_{0}\in(0,M).$ (2.12)
@@
Both the computation and the proof of Theorem 2 are easier than for the estimator minimizing (2.8), but the performance of the estimator, minimizing (1.7) is clearly inferior to that of the estimator, minimizing (1.6), see Figure 3.
@@ 3 Smooth functional theory
The smooth functional theory for the nonparametric MLE in the separated case is discussed in [piet_geurt:14] and in [piet:96] and for the non-separated case of interval censoring in [GeGr:99]. We give the treatment here for the LS estimator, minimizing (1.6), but the methods could also be used to simplify the theory given in [piet_geurt:14] and [GeGr:99] for the MLE.
We define, for the LS estimate $\hat{F}_{n}$,
$\theta_{\hat{F}_{n}}(u,v,\delta_{0},\delta_{1})$ $=\bigl\{\delta_{0}-\hat{F}_{n}(u)\bigr\}\phi_{\hat{F}_{n}}(u)+\bigl\{\delta_{1}-\hat{F}_{n}(v)+\hat{F}_{n}(u)\bigr\}\bigl\{\phi_{\hat{F}_{n}}(v)-\phi_{\hat{F}_{n}}(u)\bigr\}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad-\bigl\{\delta_{2}-(1-\hat{F}_{n}(v)\}\phi_{\hat{F}_{n}}(v).$ (3.1)
where $\phi_{\hat{F}_{n}}$ is either an integrated score function
$\phi_{\hat{F}_{n}}(x)=\int_{y\in[0,x]}a(y)\,d\hat{F}_{n}(y),$ (3.2)
@@
If we want to estimate a smooth functional of the type $\int\kappa_{F_{0}}(x)\,dF_{0}(x)$, we approximately want to have for the inner part between brackets of this integral:
@@
where the $t_{i}$ are the points of jump of $\hat{F}_{n}$ and $y_{i}=\phi_{\hat{F}_{n}}(t_{i})$; $\phi_{\hat{F}_{n}}(t)=0$ is zero if $\hat{F}_{n}(t)=0$ or $\hat{F}_{n}(t)=1$. This is analogous to (iii) of Theorem 3.1 on p. 646 of [GeGr:99]. Here $\Delta_{i}(h_{j})$ and $\Delta_{ij}(h)$ are defined by
@@
$=1.$
As an example, for the situation that $F_{0}$ is the uniform distribution function on $[0,1]$ and $(U,V)$ is uniformly distributed on the upper triangle of the unit square with vertices $(0,0)$, $(0,1)$ and $(1,1)$ and, moreover, $\kappa_{F_{0}}$ is the mean functional, this gives the solution $\phi_{\hat{F}_{n}}\equiv 1/2$. Note that this $\phi_{\hat{F}_{n}}$ does not have an interpretation as an integrated score function.
@@
Other smooth functionals can be treated in a similar way, also giving the $\sqrt{n}$ behavior and asymptotic normality of these functionals of the LS estimator, which means that we can follow the analysis in [piet:24] to show that the “off-diagonal” terms in Section 5.2
@@
where $\phi_{\hat{F}_{n}}$ is defined by (3.2), but now for the nonparametric MLE $\hat{F}_{n}$. By Proposition 3.1 on p. 644 of [GeGr:99] we have:
@@
As noted in [GeGr:99], p. 669, this is an equation of type $A{\bm{y}}={\bm{b}}$, where $A$ is a symmetrix, strictly diagonally dominant $M$-matrix (also called a Stieltjes matrix), which means that the equation has a unique solution, see [Berman_Plemmons:79]. This implies that the right-continuous piecewise constant function $\phi_{\hat{F}_{n}}$, constant on the same intervals as $\hat{F}_{n}$, is uniquely determined. Of course the same considerations hold for the equations (3).
@@
where $t_{1}<\dots<t_{m}$ are the points of jump of $F$, and $t_{0}=0$ and $t_{m+1}=1$. For this situation the information lower bound of the asymptotic variance was numerically computed to be approximately $0.1198987$, see Section 4.2 of [GeGr:99]. There is no explicit expression for it since it depends on the solution of an integral equation, for which there is no explicit form either. Based on $10,000$ simulations for sample size $1000$, we got as estimates of the asymptotic values of $n\text{var}(m(\hat{F}_{n}))$ the value $0.1237086$ for the nonparametric MLE, and the values $0.1240233$ and $0.1271471$ for the least squares estimators, minimizing (1.6) and (1.7), respectively.

### v5

@@ Abstract
The limit distribution of the nonparametric maximum likelihood estimator for interval censored data with more than one observation time per unobservable observation, is still unknown in general. For the so-called separated case, where one has observation times which are at a distance larger than a fixed epsilon¿0, the limit distribution was derived in [5]. For the non-separated case there is a conjectured limit distribution, given in [10], Section 5.2 of Part 2. Whether this conjecture holds is still unknown, but the present paper shows that for sample sizes 1000 and 10,000 this limit behavior is still not clearly seen.
We prove consistency of a related nonparametric isotonic least squares estimator and sketch of the proof for its limit distribution. We also provide simulation results to show how the nonparametric MLE and least squares estimator behave in comparison. Moreover, we discuss a simpler least squares estimator that can be computed in one step, but is inferior to the other least squares estimator, since it does not use all information.
For the simplest model of interval censoring, the current status model, the nonparametric maximum likelihood and least squares estimators are the same. This equivalence breaks down if there are more observation times per unobservable observation. The computations for the simulation of the more complicated interval censoring model were performed by using the iterative convex minorant algorithm. They are provided in the GitHub repository [7].
@@ 1 Introduction
The simplest and most studied interval censoring model is the so-called interval censoring, case 1, or current status model. This model can be defined in the following way (see, e.g., Section 2.3 of [7]).
@@
Since distribution functions are by definition nondecreasing, computing the maximum likelihood estimator poses a shape constrained optimization problem. As can be seen from (1.2), the value of $\ell$ only depends on the values that $F$ takes at the observed time points $T_{i}$. Hence one can choose to consider only distribution functions that are constant between successive observed time points $T_{i}$. The lemma below shows that this estimator can be characterized in terms of a greatest convex minorant of a certain diagram of points. The following result is Lemma 2.7 in [7] (and Proposition 1.2 in [4]).
@@
[Characterization of the nonparametric ML estimator in the current status model] Consider the cumulative sum (cusum) diagram consisting of the points $P_{0}=(0,0)$ and
@@
over all monotone nondecreasing (not necessarily bounded by 0 and $1$) functions $F$, see p. 43 in section 1.1 of part 2 of [4]. It also follows from Theorem 1.2.1 on p. 7 of [11], where the connection brtween the derivative of the greatest convex minorant and the solution of the isotonic regression problem is given.
So we have two characterizations of the nonparametric maximum likelihood estimator, the characterization as a maximizer of (1.2) and the characterization as a minimizer of (1.3). Note that the weights are equal and constant in the least sauares problem.
@@
where $\Delta_{i2}=1-\Delta_{i0}-\Delta_{i1}$. It is not clear that there is an equivalent to the minimization of (1.3) in this situation. We cannot use Theorem 1.5.1 in [11] because we have the difference of $F(V_{i})$ and $F(U_{i})$ in the terms $\log(F(V_{i})-F(U_{i}))$ instead of just $F(U_{i})$ or $F(V_{i})$ by itself only.
@@
Figure 2: (a) Simulated variances, times $n^{2/3}$, of the nonparametric MLE (black solid curve) and the least squares estimate minimizing (1.6) (red)), for $t_{i}=0.1,0.2,\dots,1.9$, linearly interpolated between values at the $t_{i}$ for the model of Example 1. The blue dashed curve is the theoretical limit curve one obtains from Theorem 1 in Section 2 for the LS estimator. The simulated variances are based on $10,000$ simulations of samples of size $n=1000$ for the truncated exponential distribution function $F_{0}$ on $[0,2]$ and the order statistics of the uniform distribution on $0,2]^{2}$ as observation times. (b) The same comparison, but now for $F_{0}$ uniform on $[0,2]$.
@@
Inspired by the conjectured faster rate of convergence of the MLE in the non-separated case (i.e., observation intervals can be arbitrarily small) in Section 5.2 of Part 2 of [4], Lucien Birgé constructed in [2] a histogram estimator which actually achieves rate $(n\log n)^{1/3}$ locally (but suffers severely from bias, in contrast with the MLE). As noted above, we still do not know whether the MLE also achieves this faster rate.
It is also possible to define a least squares estimator which can be computed in one step, using the convex minorant (or “pool adjacent violators”) algorithm. An estimator of this type was proposed for the uniform deconvolution problem in [12]. This estimator minimizes
@@
We compare the two least squares estimators for sample size $n=10,000$ in Figure 3, from which the superiority of the estimator, based on minimizing (1.6), seems clear. This second least squares estimator is in fact very much of the same type as the LS estimator/MLE in the current status model; it only checks whether the unobservable observation is larger than or smalle than a real observation time.
@@
Our proof of the limit result Theorem 1 is not complete. We still have to show that the “off-diagonal terms” (5.9) and (5.10) are $o_{p}(n^{-2/3})$, see Section 5.3. In fact, terms of this type were shown to be $O_{p}(n^{-5/6})$ in a similar problem, analyzed in [8]. In Section 3 we discuss the theory, needed for the completion of the proof.
The proof of Theorem 2 on the limit behavior of the second LS estimator is complete, and in fact very much like the proof for the limit behavior of the MLE in the current status model.
Both LS estimators lead to $\sqrt{n}$-consistent estimators of smooth functionals, but the asymptotic variance of the MLE attains the information lower bound, as proved in [3], so cannot be beaten here. The influence function of the LS estimators is in fact different from the efficient influence function of the MLE, which can be interpreted as a conditional expectation, see Section 3.
@@ 2 The least squares estimators
The proof is based on the so-called Fenchel duality condition, as also used in [7] in the characterization of the nonparametric MLE for interval censoring (apart from the Lagrange multipliers in the present case) and is therefore omitted. A picture of the process $W_{n,\hat{F}_{n}}$ for $n=100$ is given in Figure 4 for Example 1 above. $W_{n,\hat{F}_{n}}$ touches zero at points just to the left of points where the least squares estimator $\hat{F}_{n}$ has a jump. The iterative convex minorant algorithm for computing $\hat{F}_{n}$ is based on this characterization.
@@
Let $F_{0}$ have a continuous strictly positive density $f_{0}$ on $[0,M]$, where $f_{0}$ stays away from zero, and let $(U,V)$ be the order statistics of an absolutely continuous distribution. We assume that $(U,V)$ has a a positive continuous density $h$ on
@@
staying away from zero, with first and second marginals $h_{1}$ and $h_{2}$, respectively, and with bounded partial derivatives. We assume that $X_{i}$ is independent of $(U_{i},V_{i})$.
@@
$\sigma_{t_{0}}=\left(a_{t_{0}}f_{0}(t_{0})/b_{t_{0}}\right)^{2/3},\qquad t_{0}\in(0,M).$ (2.6)
@@
The conditions on the observation density $h$ are similar to the conditions in [3] and used in the smooth functional theory needed in treating so-called “off-diagonal terms”, see Section 5.3.
Lemma 3.
Let $\hat{F}_{n}$ be the isotonic least squares estimator, minimizing (1.6) over distribution functions $F$ under the conditions of Theorem 1. Then $\hat{F}_{n}$ converges almost surely to $F_{0}$ in the supremum metric.
Proof.
We use “Jewell’s method” ([10]). Let the function $\psi$ be defined by
$\psi(F)$
Since $\hat{F}_{n}$ minimizes (2) over $F$, we must have
$\geq 0.$
Proceeding as in Section 4 of Part 2 of [4], using the Helly compactness theorem, we get from this, for a limit point $F$ of a subsequence of $\hat{F}_{n}$, where $h$ is the density of $(U,V)$ (see Theorem 1):
$\int\{F_{0}(u)-F(u)\}\{F(u)-F_{0}(u)\}\,h(u,v)\,du\,dv$
$\qquad+\int\{F_{0}(v)-F_{0}(u)-F_{(}v)-F(v)+F(u)\}$
$\qquad+\int\{F_{0}(v)-F(v)\}\{F(v)-F_{0}(v)\}\,h(u,v)\,du\,dv$
$\qquad-\int\{F(v)-F_{0}(v)\}^{2}\,h(u,v)\,du\,dv$
$\geq 0.$
@@
As an example, the asymptotic variance of the LS estimator is shown in Figure 5 as the blue dashed curve for in Example 1. Remarkably, the curve of the variances of the MLE is even closer to the theoretical asymptotic curve than the curve for the variances of the least squares estimator for this sampe size of $n=10,000$. This should be compared with Figure 2, where the sample size was $n=1000$ and the variances for the LS estimator were actually smaller that those of the nonparametric MLE. The faster decrease of the variances of the MLE than those of the LS when we increase the sample size from $1000$ to $10,000$ might give a hint of the possibility of a rate of $(n\log n)^{1/3}$ again, but the form of the curve is certainly not in accordance with the conjecture in [4].
We here make some additional remarks about the proof of Theorem 1, discussed in the Appendix. Results of this type were proved for the nonparametric MLE in [6] and [8]. One has to show that the leading asymptotic behavior is provided by replacing the estimator by the underlying distribution function $F_{0}$ in the characterization, given in Lemma 2. To this end, one has to show that certain terms, called the “off-diagonal terms”, are of lower order. Doing this is rather hard in our experience ([8] needed 53 pages), but presently there seems to be no other way. We give a sketch of these calculations in Section 3 and in the Appendix. For similar calculations, see, e.g., Chapter 10 of [7] and [8]. The first result of the type of Theorem 1 was for the current status model in Theorem 5.5 of Part 2 of [4].
@@
So, in constructing the cusum diagram, we run through all $2n$ observation points; if we meet a point $U_{i}$, we add $\delta_{i0}$ to the cusum diagram, if we meet a point $V_{i}$ we add $\delta_{i0}+\delta_{i1}$ to the diagram. In the latter case we only record whether the unobservable variable lies to the left of $V_{i}$ (then $\delta_{i0}+\delta_{i1}=1$) or to the right of $V_{i}$ (then $\delta_{i0}+\delta_{i1}=0$), so we do not use the information it lies between $U_{i}$ and $V_{i}$ if this is the case.
@@
$a_{t_{0}}^{\prime}$ $=\left\{F_{0}(t_{0})\{1-F_{0}(t_{0})\}\bigl\{h_{1}(t_{0})+h_{2}(t_{0})\bigr\}\right\}^{1/2},$ (2.10)
@@
$b_{t_{0}}^{\prime}$ $=\left\{h_{1}(t_{0})+h_{2}(t_{0})\right\}/2,$ (2.11)
@@
$\sigma_{t_{0}}^{\prime}=\left(a_{t_{0}}^{\prime}f_{0}(t_{0})/b_{t_{0}}^{\prime}\right)^{2/3},\qquad t_{0}\in(0,M).$ (2.12)
@@
The proof is given in the Appendix. Both the computation and the proof of Theorem 2 are easier than for the estimator minimizing (2.7), but the performance of the estimator, minimizing (1.7) is clearly inferior to that of the estimator, minimizing (1.6), see Figure 3.
It is instructive to compare the results of Theorems 1 and 2 with the limit result for the MLE (which is also an LS estimator with equal constant weights) for the current status model. The result is given as Theorem 3.7 in [7] on p. 68. In that case we have
$n^{1/3}\left\{\hat{F}_{n}(t_{0})-F_{0}(t_{0})\right\}/\sigma\stackrel{{\scriptstyle{\cal D}}}{{\longrightarrow}}Z,$
where $Z=\text{argmin}_{t}\{W(t)+t^{2}\}$ and $\sigma=(a/b)^{2/3}$, with
$a=\left\{F_{0}(t_{0})\{1-F_{0}(t_{0})\}g(t_{0})f_{0}(t_{0})\right\}^{1/2}\qquad\text{ and }\qquad b=\tfrac{1}{2}g(t_{0}).$
Theorem 2 is completely similar, with $g(t_{0})$ replaced by $h_{1}(t_{0})+h_{2}(t_{0})$ and Theorem 1 also has this structure, but in that case $g(t_{0})$ is replaced by $2\{h_{1}(t_{0})+h_{2}(t_{0})\}$ and $a$ is replaced by a more complicated scale factor, reflecting the covariances of the trinomial distribution of $\Delta_{i0},\,\Delta_{i1}$ and $\Delta_{i2}$.
Figure 6: (a) Simulated variances, times $n^{2/3}$, of the LS estimator, minimizing (1.6) (red curve) and the LS estimator minimizing (2.7) (blue), for $t_{i}=0.1,0.2,\dots,1.9$, linearly interpolated between values at the $t_{i}$ for the model of Example 1. The simulated variances are based on $10,000$ simulations of samples of size $n=1000$ for the truncated exponential distribution function $F_{0}$ on $[0,2]$ and the order statistics of the uniform distribution on $0,2]^{2}$ as observation times. (b) The same comparison, but now for $F_{0}$ uniform on $[0,2]$.
We chose equal weights, because the MLE for current status data is equivalent to an LS estimator with equal weights. But it is also possible to introduce other weights for the LS estimators. We experimented a bit with weights of the form $w_{i0}=1/U_{i}$, $w_{i1}=1/(V_{i}-U_{i})$ and $w_{i2}=1/(2-V_{i})$, $i=1,\dots,n$, for our examples of the truncated exponential distribution and the Uniform distribution on $[0,2]$, giving smaller intervals more weight. Our findings were that the iterative procedure for computing the LS estimator, minimizing
$\sum_{i=1}^{n}\left\{\left\{F(U_{i})-\Delta_{i0}\right\}^{2}w_{i0}+\left\{F(V_{i})-F(U_{i})-\Delta_{i1}\right\}^{2}w_{i1}+\left\{1-F(V_{i})-\Delta_{i2}\right\}^{2}w_{i2}\right\}$ (2.14)
lasted longer and became less stable. Also, the results for the algorithm with equal weights were slightly better, for sample size $n=1000$, see Figure 6.
@@ 3 Smooth functional theory
The smooth functional theory for the nonparametric MLE in the separated case is discussed in [7] and in [6] and for the non-separated case of interval censoring in [3]. We give the treatment here for the LS estimator, minimizing (1.6). We define, for a distribution function $F$, the function $\theta_{F}$ by:
$\theta_{F}(u,v,\delta_{0},\delta_{1})$ $=\bigl\{\delta_{0}-F(u)\bigr\}\phi_{F}(u)+\bigl\{\delta_{1}-F(v)+F(u)\bigr\}\bigl\{\phi_{F}(v)-\phi_{F}(u)\bigr\}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad-\bigl\{\delta_{2}-(1-F(v)\}\phi_{\hat{F}_{n}}(v).$ (3.1)
where $\phi_{F}$ is either an integrated score function
$\phi_{F}(x)=\int_{y\in[0,x]}a(y)\,dF(y),$ (3.2)
@@
We also have:
$\int\theta_{F_{0}}\,dQ_{0}=0,$
for the underlying probability measure $Q_{0}$.
Note that $\theta_{F}$ is similar to, but different from, the canonical gradient or “efficient influence function” $\tilde{\theta}=\tilde{\theta}_{F,H}$ in [3] (see, e.g., the discussion preceding Theorem 2.1 in [3]). The theory in [3] shows that the information lower bound $\|\tilde{\theta}\|^{2}_{Q_{0}}$, attained by the asymptotic variance of the MLE in estimating smooth functionals, cannot be improved upon, so we can only hope that the corresponding asymptotic variance $\|\theta_{F_{0}}\|^{2}_{Q_{0}}$ of the LS estimator (see below) will be close to this information lower bound.
@@
If we want to estimate a smooth functional of the type $\int\kappa_{F_{0}}(x)\,dF_{0}(x)$, we approximately want to accomplish for the inner part between brackets of this integral:
@@
where the $t_{i}$ are the points of jump of $\hat{F}_{n}$ and $y_{i}=\phi_{\hat{F}_{n}}(t_{i})$; $\phi_{\hat{F}_{n}}(t)=0$ is zero if $\hat{F}_{n}(t)=0$ or $\hat{F}_{n}(t)=1$. This is analogous to (iii) of Theorem 3.1 on p. 646 of [3]. Here $\Delta_{i}(h_{j})$ and $\Delta_{ij}(h)$ are defined by
@@
$=1,\qquad x\in\{z\in(0,M):0<\hat{F}_{n}(z)<1\}$
As an example, for the situation that $F_{0}$ is the uniform distribution function on $[0,1]$ and $(U,V)$ is uniformly distributed on the upper triangle of the unit square with vertices $(0,0)$, $(0,1)$ and $(1,1)$ and, moreover, $\kappa_{F_{0}}$ is the mean functional, this gives the solution
$\phi_{\hat{F}_{n}}(x)=\left\{\begin{array}[]{ll}1/2,\qquad&x\in\{z\in(0,M):0<\hat{F}_{n}(z)<1\}\\ 0,&\text{otherwise}.\end{array}\right$.
@@
Other smooth functionals can be treated in a similar way, also giving the $\sqrt{n}$ behavior and asymptotic normality of these functionals of the LS estimator, which means that we can follow the analysis in [8] to show that the “off-diagonal” terms in Section 5.3
@@
where $\phi_{\hat{F}_{n}}$ is defined by (3.2), but now for the nonparametric MLE $\hat{F}_{n}$. By Proposition 3.1 on p. 644 of [3] we have:
@@
As noted in [3], p. 669, this is an equation of type $A{\bm{y}}={\bm{b}}$, where $A$ is a symmetrix, strictly diagonally dominant $M$-matrix (also called a Stieltjes matrix), which means that the equation has a unique solution, see [1]. This implies that the right-continuous piecewise constant function $\phi_{\hat{F}_{n}}$, constant on the same intervals as $\hat{F}_{n}$, is uniquely determined. Of course the same considerations hold for the equations (3).
@@
where $t_{1}<\dots<t_{m}$ are the points of jump of $F$, and $t_{0}=0$ and $t_{m+1}=1$. For this situation the information lower bound of the asymptotic variance was numerically computed to be approximately $0.1198987$, see Section 4.2 of [3]. There is no explicit expression for it since it depends on the solution of an integral equation, for which there is no explicit form either. Based on $10,000$ simulations for sample size $1000$, we got as estimates of the asymptotic values of $n\text{var}(m(\hat{F}_{n}))$ the value $0.1237086$ for the nonparametric MLE, and the values $0.1240233$ and $0.1271471$ for the least squares estimators, minimizing (1.6) and (1.7), respectively.